OpenStudy (aroub):

Given that 20m=50t find m+t/m-t

OpenStudy (anonymous):

it's \( \frac{m+t}{m-t} \) isn't ?

OpenStudy (aroub):

yes

OpenStudy (anonymous):

m =2.5t m+t/m-t = 3.5t / 0.5t = 7

OpenStudy (anonymous):

\( 20m = 50t \Rightarrow m = \frac 52 t \) now substitute in \( \frac{m+t}{m-t} \)... can you do the rest ?

OpenStudy (aroub):

Oh and btw the answer is 7/3 ( i just don't know how to get there :( @Fool , I just substitute the m and solve?

OpenStudy (anonymous):

yes exactly.

hero (hero):

It's not that hard

OpenStudy (anonymous):

and you will have 7/3! :)

OpenStudy (anonymous):

definitely not for a Hero.

OpenStudy (anonymous):

where did i go wrong?? oh yes - i cant subtract!!!!!

OpenStudy (aroub):

jimmy, you substituted the m ? like fool said?

hero (hero):

Hmm, I might have to participate in this

OpenStudy (aroub):

okay...

hero (hero):

How come FFM didn't show his steps

OpenStudy (aroub):

FFM?

hero (hero):

Fool for math

OpenStudy (aroub):

oh

hero (hero):

I got it

hero (hero):

You have to substitute twice in order to get 7/3

hero (hero):

Or just put everything in terms of one variable

hero (hero):

Aroub, why did you disappear?

OpenStudy (aroub):

ok can you show me the steps please?

hero (hero):

Yes, hold on a minute

OpenStudy (aroub):

ok thank you :)

hero (hero):

http://vyew.com/room#/418625/Solved

hero (hero):

Go there

OpenStudy (anonymous):